# Provable Benefit of Curriculum in Transformer Tree-Reasoning Post-Training

### v1

@@ D.2 Proof of Finetuning Algorithms
$\nabla_{\mathbf{W}^{k^{\star}}}\mathcal{J}_{\mathrm{REINFORCE}}^{k^{\star}}$ $\stackrel{{\scriptstyle(1)}}{{=}}\nabla_{\mathbf{W}^{k^{\star}}}\int R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\,p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})\,d\tau$ (26)
$\stackrel{{\scriptstyle(2)}}{{=}}\int R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\,\nabla_{\mathbf{W}^{k^{\star}}}p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})\,d\tau$ (27)
$\stackrel{{\scriptstyle(3)}}{{=}}\int R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\,p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})\,\nabla_{\mathbf{W}^{k^{\star}}}\log p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})\,d\tau$ (28)
$\stackrel{{\scriptstyle(4)}}{{=}}\mathbb{E}_{\tau}\Big[R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\sum_{l=1}^{k^{\star}+1}\nabla_{\mathbf{W}^{k^{\star}}}\log\pi_{\mathbf{W}^{k^{\star}}}\big({i}_{l}\mid\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}}\big)\Big],$ (29)
@@
Finally, we note that deterministic $\phi$ does not affect the score terms and only enters through $R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)$, preserving all interchanges above. Fubini’s theorem applies whenever $\mathbb{E}_{\tau}[\,|R^{k^{\star}}\sum_{l}\nabla\log\pi_{l}|\,]<\infty$, which holds by boundedness of $R^{k^{\star}}$ and square-integrability (log-smoothness) of the score.
@@
$\alpha_{l}^{(t)}(j)=\mathrm{softmax}_{j}\big(\mathbf{p}_{j}^{\top}\mathbf{W}^{(t)}\mathbf{p}_{c_{l}}+m_{l}(j)\big),\quad\eta_{l}^{(t)}(j)=\frac{\langle\bm{\mu}_{x_{j}}-\hat{\bm{\mu}}^{\mathrm{att},(t)}_{l+1},\ \bm{\mu}_{x_{i_{l}}}-\sum_{r}p^{(t)}_{\mathrm{vocab}}(r)\bm{\mu}_{r}\rangle}{\beta}.$
@@
$p_{\mathrm{path}}\ :=\ \mathbb{E}_{\mathbf{x}}\,\mathbb{P}_{\tau\mid\mathbf{x}}\big\{\tau\in\mathcal{T}_{\mathrm{true}}\big\}=\mathbb{E}_{\mathbf{x}}\Big[\prod_{l=1}^{k^{\star}+1}\pi_{\mathbf{W}^{k^{\star}}}\big(i_{l}\mid\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}}\big)\Big].$
@@
$\rho_{\mathrm{spur}}\ :=\ \mathbb{E}_{\mathbf{x}}\Big[\mathbb{P}_{\tau\mid\mathbf{x}}\big\{R^{k^{\star}}(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau))=1\ \big|\ \tau\notin\mathcal{T}_{\mathrm{true}}\big\}\Big],$
@@
$\delta\hat{\bm{\mu}}^{\mathrm{att},(t)}_{l+1}\ =\ \sum_{k\in\mathcal{I}_{l}}\delta\alpha_{l}^{(t)}(k)\,\bm{\mu}_{x_{k}}\ +\ \mathcal{O}(\eta^{2}),$ (39)
$\delta p^{(t)}_{\mathrm{vocab}}(r)\ =\ p^{(t)}_{\mathrm{vocab}}(r)\,\frac{\big\langle\delta\hat{\bm{\mu}}^{\mathrm{att},(t)}_{l+1},\ \bm{\mu}_{r}-\sum_{q}p^{(t)}_{\mathrm{vocab}}(q)\bm{\mu}_{q}\big\rangle}{\beta}\ +\ \mathcal{O}(\eta^{2}),$ (40)
@@
(ii) Attention-weighted token vector $\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}$. Since $\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}=\sum_{j}\alpha_{l}(j)\,\bm{\mu}_{x_{j}}$ is linear in $\bm{\alpha}_{l}$,
$\delta\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}=\sum_{j}\delta\alpha_{l}(j)\,\bm{\mu}_{x_{j}}=\sum_{j,k}J^{\alpha}_{jk}\,\delta s_{l}(k)\,\bm{\mu}_{x_{j}}+\frac{1}{2}\sum_{j,k,m}H^{\alpha}_{j,km}\,\delta s_{l}(k)\,\delta s_{l}(m)\,\bm{\mu}_{x_{j}}+\mathcal{O}(\eta^{3}),$
so $\delta\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}=\mathcal{O}(\eta)+\mathcal{O}(\eta^{2})$.
(iii) Vocabulary softmax $p_{\mathrm{vocab}}$. Let $\ell_{r}=\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{r}\rangle/\beta$ and $\bm{p}=\mathrm{softmax}(\bm{\ell})$. Then
$\delta\ell_{r}=\frac{\langle\delta\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\ \bm{\mu}_{r}\rangle}{\beta},\qquad\delta p(r)=\sum_{m}J^{p}_{rm}\,\delta\ell_{m}+\frac{1}{2}\sum_{m,n}H^{p}_{r,mn}\,\delta\ell_{m}\,\delta\ell_{n}+\mathcal{O}(\eta^{3}),$
@@
Since $\delta\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}=\mathcal{O}(\eta)+\mathcal{O}(\eta^{2})$, we have $\delta\ell=\mathcal{O}(\eta)+\mathcal{O}(\eta^{2})$ and thus $\delta p=\mathcal{O}(\eta)+\mathcal{O}(\eta^{2})$.
@@
$\eta_{l}(k)=\frac{1}{\beta}\Big\langle\underbrace{\bm{\mu}_{x_{k}}-\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}}_{\displaystyle:=\,\mathbf{u}},\ \underbrace{\bm{\mu}_{x_{i_{l}}}-\sum_{r}p(r)\,\bm{\mu}_{r}}_{\displaystyle:=\,\mathbf{v}}\Big\rangle.$
Perturbing $(\mathbf{u},\mathbf{v})\mapsto(\mathbf{u}-\delta\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\ \mathbf{v}-\sum_{r}\delta p(r)\,\bm{\mu}_{r})$ and expanding to second order yields
$\delta\eta_{l}(k)=\frac{1}{\beta}\Big(-\langle\delta\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\ \mathbf{v}\rangle-\big\langle\mathbf{u},\ \sum_{r}\delta p(r)\,\bm{\mu}_{r}\big\rangle\Big)+\frac{1}{\beta}\Big(-\tfrac{1}{2}\langle\delta^{2}\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\ \mathbf{v}\rangle-\big\langle\delta\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\ \sum_{r}\delta p(r)\,\bm{\mu}_{r}\big\rangle-\tfrac{1}{2}\big\langle\mathbf{u},\ \sum_{r}\delta^{2}p(r)\,\bm{\mu}_{r}\big\rangle\Big)+\mathcal{O}(\eta^{3}),$
where $\delta^{2}\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}$ and $\delta^{2}p(r)$ collect the quadratic terms shown in (ii) and (iii). Since $\delta\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}=\mathcal{O}(\eta)$ and $\delta p=\mathcal{O}(\eta)$ at leading order, all bracketed second-line contributions are $\mathcal{O}(\eta^{2})$.
@@
$\begin{aligned} &\widehat{\mathcal{J}}^{k^{\star}}_{n}(\mathbf{W}^{k^{\star}}):=\frac{1}{n}\sum_{s=1}^{n}R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau^{(s)}_{-2})\big),\\ &\widehat{\mathbf{g}}^{k^{\star}}_{n}(\mathbf{W}^{k^{\star}}):=\frac{1}{n}\sum_{s=1}^{n}R^{k^{\star}}(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau^{(s)}_{-2}))\,\sum_{l=1}^{k^{\star}+1}\nabla_{\mathbf{W}^{k^{\star}}}\log\pi_{\mathbf{W}^{k^{\star}}}\big({i}^{(s)}_{l}\mid\mathbf{x}^{(s)},\hat{\bm{\mu}}^{z_{1:l}}\big).\end{aligned}$
@@
$p_{\mathrm{succ}}(\mathbf{W}^{k^{\star}}):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}_{x}}\Big[\mathbb{P}_{\tau\sim p_{\mathbf{W}^{k^{\star}}}(\cdot\mid\mathbf{x})}\big\{R^{k^{\star}}(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2}))=1\big\}\Big]=\mathbb{E}_{\mathbf{x},\tau}\big[R^{k^{\star}}(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2}))\big].$
@@
where $\pi_{l}$ abbreviates $\pi_{\mathbf{W}^{k^{\star}}}({i}_{l}\mid\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}})$ and $\otimes$ is the outer product in parameter space.
@@
$\mathbf{G}\ :=\ R^{k^{\star}}\sum_{l=1}^{k^{\star}+1}\nabla_{\mathbf{W}^{k^{\star}}}\log\pi_{\mathbf{W}^{k^{\star}}}\big({i}_{l}\mid\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}}\big).$
@@
$\big|\eta_{l}^{k^{\star}}(k)\big|=\frac{\big|\langle\bm{\mu}_{x_{k}}-\hat{\bm{\mu}}^{\mathrm{att}},\ \bm{\mu}_{x_{i_{l}}}-\sum_{r}p_{\mathrm{vocab}}(r)\bm{\mu}_{r}\rangle\big|}{\beta}\ \leq\ \frac{\|\bm{\mu}_{x_{k}}-\hat{\bm{\mu}}^{\mathrm{att}}\|_{2}\,\|\bm{\mu}_{x_{i_{l}}}-\sum_{r}p_{\mathrm{vocab}}(r)\bm{\mu}_{r}\|_{2}}{\beta}\ \leq\ \frac{4}{\beta},$
@@
5:  $\widehat{\mathbf{g}}=\frac{1}{n_{0}}\sum_{s=1}^{n_{0}}\mathbf{R}^{f_{S_{\star}}}_{\mathbf{x}^{(s)}}(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau^{(s)}_{-2}))\sum_{\ell=1}^{k^{\star}+1}\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}^{(0)}}({i}^{(s)}_{\ell}\mid\mathbf{x}^{(s)},\hat{\bm{\mu}}^{z_{1:\ell}})$.
@@
5:  For each sample $s$, truncate trajectory at length $\ell+1$:
@@
8:  Compute empirical REINFORCE gradient for step $\ell$:
9:   $\widehat{\mathbf{g}}_{\ell}=\frac{1}{n_{\ell}}\sum_{s=1}^{n_{\ell}}\mathbf{R}^{\mathcal{F}_{S_{\star}}}_{\mathbf{x}^{(s)}}(\tau^{(s)}_{\text{trunc}})\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}^{(\ell-1)}}({i}^{(s)}_{\ell}\mid\mathbf{x}^{(s)},\hat{\bm{\mu}}^{z_{1:\ell}})$.
10:  Update parameters: $\mathbf{W}^{(\ell)}=\mathbf{W}^{(\ell-1)}+\eta_{\ell}^{\mathsf{B}}\cdot\widehat{\mathbf{g}}_{\ell}$.
11:end for
12: Return $\mathbf{W}^{(k^{\star}+1)}$.
@@
9:   $\widehat{\mathbf{g}}_{\ell}=\frac{1}{n_{\ell}}\sum_{s=1}^{n_{\ell}}\mathbf{R}^{\mathcal{F}_{S_{\star}}}_{\mathbf{x}^{(s)}}(\tau^{(s)}_{\text{hint}})\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}^{(k^{\star}+1-\ell)}}({i}^{(s)}_{\ell}\mid\mathbf{x}^{(s)},\hat{\bm{\mu}}^{z_{1:\ell}})$.
@@
D.3 Proof of Test-time Scaling

### v2

@@ G.2 Proof of Finetuning Algorithms
$\nabla_{\mathbf{W}^{k^{\star}}}\mathcal{J}_{\mathrm{REINFORCE}}^{k^{\star}}$ $\stackrel{{\scriptstyle(1)}}{{=}}\nabla_{\mathbf{W}^{k^{\star}}}\int R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\,p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})\,d\tau$ (26)
$\stackrel{{\scriptstyle(2)}}{{=}}\int R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\,\nabla_{\mathbf{W}^{k^{\star}}}p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})\,d\tau$ (27)
$\stackrel{{\scriptstyle(3)}}{{=}}\int R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\,p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})\,\nabla_{\mathbf{W}^{k^{\star}}}\log p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})\,d\tau$ (28)
$\stackrel{{\scriptstyle(4)}}{{=}}\mathbb{E}_{\tau}\Big[R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\sum_{l=1}^{k^{\star}+1}\nabla_{\mathbf{W}^{k^{\star}}}\log\pi_{\mathbf{W}^{k^{\star}}}\big({i}_{l}\mid\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}}\big)\Big],$ (29)
@@
Finally, we note that deterministic $\phi$ does not affect the score terms and only enters through $R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)$, preserving all interchanges above. Fubini’s theorem applies whenever $\mathbb{E}_{\tau}[\,|R^{k^{\star}}\sum_{l}\nabla\log\pi_{l}|\,]<\infty$, which holds by boundedness of $R^{k^{\star}}$ and square-integrability (log-smoothness) of the score.
@@
$\alpha_{l}^{(t)}(j)=\mathrm{softmax}_{j}\big(\mathbf{p}_{j}^{\top}\mathbf{W}^{(t)}\mathbf{p}_{c_{l}}+m_{l}(j)\big),\quad\eta_{l}^{(t)}(j)=\frac{\langle\mathbf{p}_{j}-\hat{\mathbf{p}}^{\mathrm{att},(t)}_{l+1},\ \mathbf{p}_{{i_{l}}}-\sum_{r}p^{(t)}_{\mathrm{vocab}}(r)\mathbf{p}_{r}\rangle}{\beta}.$
@@
$p_{\mathrm{path}}\ :=\ \mathbb{E}_{\mathbf{x}}\,\mathbb{P}_{\tau\mid\mathbf{x}}\big\{\tau\in\mathcal{T}_{\mathrm{true}}\big\}=\mathbb{E}_{\mathbf{x}}\Big[\prod_{l=1}^{k^{\star}+1}\pi_{\mathbf{W}^{k^{\star}}}\big(i_{l}\mid\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}}\big)\Big].$
@@
$\rho_{\mathrm{spur}}\ :=\ \mathbb{E}_{\mathbf{x}}\Big[\mathbb{P}_{\tau\mid\mathbf{x}}\big\{R^{k^{\star}}(\hat{\mathbf{p}}^{z_{k^{\star}}}(\tau))=1\ \big|\ \tau\notin\mathcal{T}_{\mathrm{true}}\big\}\Big],$
@@
$\delta\hat{\mathbf{p}}^{\mathrm{att},(t)}_{l+1}\ =\ \sum_{k\in\mathcal{I}_{l}}\delta\alpha_{l}^{(t)}(k)\,\mathbf{p}_{k}\ +\ \mathcal{O}(\eta^{2}),$ (39)
$\delta p^{(t)}_{\mathrm{vocab}}(r)\ =\ p^{(t)}_{\mathrm{vocab}}(r)\,\frac{\big\langle\delta\hat{\mathbf{p}}^{\mathrm{att},(t)}_{l+1},\ \mathbf{p}_{r}-\sum_{q}p^{(t)}_{\mathrm{vocab}}(q)\mathbf{p}_{q}\big\rangle}{\beta}\ +\ \mathcal{O}(\eta^{2}),$ (40)
@@
(ii) Attention-weighted token vector $\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}$. Since $\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}=\sum_{j}\alpha_{l}(j)\,\mathbf{p}_{j}$ is linear in $\bm{\alpha}_{l}$,
$\delta\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}=\sum_{j}\delta\alpha_{l}(j)\,\mathbf{p}_{j}=\sum_{j,k}J^{\alpha}_{jk}\,\delta s_{l}(k)\,\mathbf{p}_{j}+\frac{1}{2}\sum_{j,k,m}H^{\alpha}_{j,km}\,\delta s_{l}(k)\,\delta s_{l}(m)\,\mathbf{p}_{j}+\mathcal{O}(\eta^{3}),$
so $\delta\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}=\mathcal{O}(\eta)+\mathcal{O}(\eta^{2})$.
(iii) Vocabulary softmax $p_{\mathrm{vocab}}$. Let $\ell_{r}=\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{r}\rangle/\beta$ and $\bm{p}=\mathrm{softmax}(\bm{\ell})$. Then
$\delta\ell_{r}=\frac{\langle\delta\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\ \mathbf{p}_{r}\rangle}{\beta},\qquad\delta p(r)=\sum_{m}J^{p}_{rm}\,\delta\ell_{m}+\frac{1}{2}\sum_{m,n}H^{p}_{r,mn}\,\delta\ell_{m}\,\delta\ell_{n}+\mathcal{O}(\eta^{3}),$
@@
Since $\delta\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}=\mathcal{O}(\eta)+\mathcal{O}(\eta^{2})$, we have $\delta\ell=\mathcal{O}(\eta)+\mathcal{O}(\eta^{2})$ and thus $\delta p=\mathcal{O}(\eta)+\mathcal{O}(\eta^{2})$.
@@
$\eta_{l}(k)=\frac{1}{\beta}\Big\langle\underbrace{\mathbf{p}_{k}-\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}}_{\displaystyle:=\,\mathbf{u}},\ \underbrace{\mathbf{p}_{{i_{l}}}-\sum_{r}p(r)\,\mathbf{p}_{r}}_{\displaystyle:=\,\mathbf{v}}\Big\rangle.$
Perturbing $(\mathbf{u},\mathbf{v})\mapsto(\mathbf{u}-\delta\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\ \mathbf{v}-\sum_{r}\delta p(r)\,\mathbf{p}_{r})$ and expanding to second order yields
$\delta\eta_{l}(k)=\frac{1}{\beta}\Big(-\langle\delta\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\ \mathbf{v}\rangle-\big\langle\mathbf{u},\ \sum_{r}\delta p(r)\,\mathbf{p}_{r}\big\rangle\Big)+\frac{1}{\beta}\Big(-\tfrac{1}{2}\langle\delta^{2}\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\ \mathbf{v}\rangle-\big\langle\delta\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\ \sum_{r}\delta p(r)\,\mathbf{p}_{r}\big\rangle-\tfrac{1}{2}\big\langle\mathbf{u},\ \sum_{r}\delta^{2}p(r)\,\mathbf{p}_{r}\big\rangle\Big)+\mathcal{O}(\eta^{3}),$
where $\delta^{2}\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}$ and $\delta^{2}p(r)$ collect the quadratic terms shown in (ii) and (iii). Since $\delta\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}=\mathcal{O}(\eta)$ and $\delta p=\mathcal{O}(\eta)$ at leading order, all bracketed second-line contributions are $\mathcal{O}(\eta^{2})$.
@@
$\begin{aligned} &\widehat{\mathcal{J}}^{k^{\star}}_{n}(\mathbf{W}^{k^{\star}}):=\frac{1}{n}\sum_{s=1}^{n}R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau^{(s)}_{-2})\big),\\ &\widehat{\mathbf{g}}^{k^{\star}}_{n}(\mathbf{W}^{k^{\star}}):=\frac{1}{n}\sum_{s=1}^{n}R^{k^{\star}}(\hat{\mathbf{p}}^{z_{k^{\star}}}(\tau^{(s)}_{-2}))\,\sum_{l=1}^{k^{\star}+1}\nabla_{\mathbf{W}^{k^{\star}}}\log\pi_{\mathbf{W}^{k^{\star}}}\big({i}^{(s)}_{l}\mid\mathbf{x}^{(s)},\hat{\mathbf{p}}^{z_{1:l}}\big).\end{aligned}$
@@
$p_{\mathrm{succ}}(\mathbf{W}^{k^{\star}}):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}_{x}}\Big[\mathbb{P}_{\tau\sim p_{\mathbf{W}^{k^{\star}}}(\cdot\mid\mathbf{x})}\big\{R^{k^{\star}}(\hat{\mathbf{p}}^{z_{k^{\star}}}(\tau_{-2}))=1\big\}\Big]=\mathbb{E}_{\mathbf{x},\tau}\big[R^{k^{\star}}(\hat{\mathbf{p}}^{z_{k^{\star}}}(\tau_{-2}))\big].$
@@
where $\pi_{l}$ abbreviates $\pi_{\mathbf{W}^{k^{\star}}}({i}_{l}\mid\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}})$ and $\otimes$ is the outer product in parameter space.
@@
$\mathbf{G}\ :=\ R^{k^{\star}}\sum_{l=1}^{k^{\star}+1}\nabla_{\mathbf{W}^{k^{\star}}}\log\pi_{\mathbf{W}^{k^{\star}}}\big({i}_{l}\mid\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}}\big).$
@@
$\big|\eta_{l}^{k^{\star}}(k)\big|=\frac{\big|\langle\mathbf{p}_{k}-\hat{\mathbf{p}}^{\mathrm{att}},\ \mathbf{p}_{{i_{l}}}-\sum_{r}p_{\mathrm{vocab}}(r)\mathbf{p}_{r}\rangle\big|}{\beta}\ \leq\ \frac{\|\mathbf{p}_{k}-\hat{\mathbf{p}}^{\mathrm{att}}\|_{2}\,\|\mathbf{p}_{{i_{l}}}-\sum_{r}p_{\mathrm{vocab}}(r)\mathbf{p}_{r}\|_{2}}{\beta}\ \leq\ \frac{4}{\beta},$
@@
5:  $\widehat{\mathbf{g}}=\frac{1}{n_{0}}\sum_{s=1}^{n_{0}}\mathbf{R}^{f_{S_{\star}}}_{\mathbf{x}^{(s)}}(\hat{\mathbf{p}}^{z_{k^{\star}}}(\tau^{(s)}_{-2}))\sum_{\ell=1}^{k^{\star}+1}\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}^{(0)}}({i}^{(s)}_{\ell}\mid\mathbf{x}^{(s)},\hat{\mathbf{p}}^{z_{1:\ell}})$.
@@
5:  If $\ell<k^{\star}+1$, for each sample $s$, truncate trajectory at length $\ell+1$:
@@
8:  If $\ell=k^{\star}+1$, $\tau^{(s)}_{\text{trunc}}=\tau^{(s)}$.
9:  Compute empirical REINFORCE gradient for step $\ell$:
10:   $\widehat{\mathbf{g}}_{\ell}=\frac{1}{n_{\ell}}\sum_{s=1}^{n_{\ell}}\mathbf{R}^{\mathcal{F}_{S_{\star}}}_{\mathbf{x}^{(s)}}(\tau^{(s)}_{\text{trunc}})\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}^{(\ell-1)}}({i}^{(s)}_{\ell}\mid\mathbf{x}^{(s)},\hat{\mathbf{p}}^{z_{1:\ell}})$.
11:  Update parameters: $\mathbf{W}^{(\ell)}=\mathbf{W}^{(\ell-1)}+\eta_{\ell}^{\mathsf{B}}\cdot\widehat{\mathbf{g}}_{\ell}$.
12:end for
13: Return $\mathbf{W}^{(k^{\star}+1)}$.
@@
9:   $\widehat{\mathbf{g}}_{\ell}=\frac{1}{n_{\ell}}\sum_{s=1}^{n_{\ell}}\mathbf{R}^{\mathcal{F}_{S_{\star}}}_{\mathbf{x}^{(s)}}(\tau^{(s)}_{\text{hint}})\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}^{(k^{\star}+1-\ell)}}({i}^{(s)}_{\ell}\mid\mathbf{x}^{(s)},\hat{\mathbf{p}}^{z_{1:\ell}})$.
@@
G.3 Proof of Test-time Scaling